# Forecasts for Galaxy Formation and Dark Matter Constraints from Dwarf Galaxy Surveys

### v1

@@ Abstract
The abundance of faint dwarf galaxies is determined by the underlying population of low-mass dark matter (DM) halos and the efficiency of galaxy formation in these systems. Here, we quantify potential galaxy formation and DM constraints from future dwarf satellite galaxy surveys. We generate satellite populations using a suite of Milky Way (MW)-mass cosmological zoom-in simulations and an empirical galaxy–halo connection model, and assess sensitivity to galaxy formation and DM signals when marginalizing over galaxy–halo connection uncertainties. We find that a survey of all satellites around one MW-mass host can constrain a galaxy formation cutoff at peak virial masses of $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$ at the $1\sigma$ level; however, a tail toward low $\mathcal{M}_{50}$ prevents a $2\sigma$ measurement. In this scenario, combining hosts with differing bright satellite abundances significantly reduces uncertainties on $\mathcal{M}_{50}$ at the $1\sigma$ level, but the $2\sigma$ tail toward low $\mathcal{M}_{50}$ persists. We project that observations of one (two) complete satellite populations can constrain warm DM models with $m_{\mathrm{WDM}}\approx 10~{}\mathrm{keV}$ ($20~{}\mathrm{keV}$). Subhalo mass function (SHMF) suppression can be constrained to $\approx 70\%$, $60\%$, and $50\%$ that in CDM at peak virial masses of $10^{8}$, $10^{9}$, and $10^{10}~{}M_{\mathrm{\odot}}$, respectively; SHMF enhancement constraints are weaker ($\approx 20$, $4$, and $2$ times that in CDM, respectively) due to galaxy–halo connection degeneracies. These results motivate searches for faint dwarf galaxies beyond the MW and indicate that ongoing missions like Euclid and upcoming facilities including the Vera C. Rubin Observatory and Nancy Grace Roman Space Telescope will probe new galaxy formation and DM physics.
@@ 1 Introduction
The faintest dwarf galaxies inhabit the smallest dark matter halos that can form galaxies. These systems are sensitive probes of galaxy formation models and simultaneously test dark matter (DM) particle physics (e.g., Bullock & Boylan-Kolchin 2017; Gluscevic et al. 2019; Simon 2019; Nadler 2021; Sales et al. 2022), including DM free-streaming, interactions, and decays, wave interference of ultra-light DM, and non-standard DM production mechanisms (e.g., see Bechtol et al. 2022 for an overview).
@@
To disentangle the information dwarf galaxy surveys can offer, it is crucial to understand the interplay between galaxy formation and DM physics in low-mass halos. For example, reionization inhibits star formation in halos with peak masses below $\approx 10^{8}$ to $10^{9}~{}M_{\mathrm{\odot}}$ because of their low virial temperatures and limited gas reservoirs (e.g., Benitez-Llambay & Frenk 2020; Munshi et al. 2021; Kravtsov & Manwadkar 2022; Ahvazi et al. 2023), suppressing the population of dwarf galaxies that occupy these systems. Observations of the faintest dwarf galaxies can shed light on these processes and can also reveal deviations in underlying halo populations from CDM, which would signal new DM physics on small scales. Moreover, a measurement of the galaxy formation cutoff is necessary to infer the existence of “dark” (galaxy–free) halos—which remain a key, untested predicted of many DM models—in combination with gravitational probes of even smaller halos (e.g., Banik et al. 2021b; Gilman et al. 2020a; Nadler et al. 2021b).
@@
Current constraints on the galaxy formation cutoff are limited by the small number of known ultra-faint dwarf galaxies, which have almost exclusively been detected as satellites of the Milky Way (MW; e.g., Drlica-Wagner et al. 2020). Additionally, this measurement is plagued by large theoretical uncertainties on the faint end of the galaxy–halo connection (e.g., see Wechsler & Tinker 2018 for a review). Recently, Nadler et al. (2020b) analyzed the full-sky MW satellite population detected by the Dark Energy Survey (DES; Abbott et al. 2018) and Pan-STARRS1 (PS1; Chambers et al. 2016) to place a $95\%$ confidence upper limit of $8.5\times 10^{7}~{}M_{\mathrm{\odot}}$ on the peak virial mass at which $50\%$ of halos host galaxies, hereafter $\mathcal{M}_{50}$. [^1] We define “galaxies” as dark matter-dominated systems with absolute $V$-band magnitudes $M_{V}<0~{}\rm{mag}$, which are composed of $\gtrsim 100$ stars in our fiducial models, on average, following Nadler et al. (2020b). This constraint accounts for observational selection effects, marginalizes over the relevant theoretical uncertainties, and implies stringent limits on a variety of DM scenarios that suppress the abundance of halos with peak masses of $\approx 10^{8}~{}M_{\mathrm{\odot}}$ (e.g., Nadler et al. 2021c; Newton et al. 2021; Dekker et al. 2022). However, current data do not permit a measurement of the galaxy formation cutoff because only tens of ultra-faints have been detected; this number is too small to rule out statistical fluctuations that mimic a cutoff or to overcome theoretical uncertainties on the predicted abundance of MW satellite galaxies.
@@
Upcoming observational facilities will substantially improve this situation by expanding our census of faint dwarf galaxies. For example, the Vera C. Rubin Observatory (Rubin; Ivezić et al. 2008) is expected to detect much of the remaining MW satellite population, as well as many resolved dwarf galaxies out to $\approx 5~{}\mathrm{Mpc}$ (Drlica-Wagner et al., 2019; Mutlu-Pakdil et al., 2021); comparable sensitivity has been projected for the Chinese Space Station Telescope (CSST; Qu et al. 2023). The Nancy Grace Roman Space Telescope (Roman; Spergel et al. 2015) will also be a powerful dwarf galaxy discovery machine (Gezari et al., 2022), with the ability to detect resolved stars out to even larger distances than Rubin. The Euclid Dark Energy Mission (Euclid Collaboration et al. 2022; launched 1 July 2023) is expected to achieve similar sensitivity (Racca et al., 2016). Meanwhile, despite its small field of view, JWST can resolve stars out to much larger distances than Roman or Euclid and thus may play an important role in detecting faint dwarf galaxies throughout the Local Volume and beyond (e.g., Conroy et al. 2023). Looking further ahead, next-generation optical/infrared space telescopes such as the Habitable Worlds Observatory, which is the top priority of the Astro2020 Decadal Survey, will revolutionize the search for dwarf galaxies at cosmic distances (e.g., see Chapter 5 of The LUVOIR Team (2019) for estimates of future large space telescopes’ detection sensitivities).
@@
To prepare for these advances, we study measurements of galaxy formation and DM physics in the context of future dwarf galaxy surveys. In particular, we use a suite of MW-mass cosmological zoom-in simulations (Nadler et al., 2023) to assess the sensitivity of satellite galaxy populations to a reionization-like galaxy formation cutoff, a suppression of low-mass halo abundances characteristic of warm dark matter (WDM), and the amplitude of the subhalo mass function (SHMF). We focus on subhalos and satellite galaxies (rather than isolated halos and field galaxies) because, as discussed above, future observational facilities capable of resolving ultra-faint dwarfs beyond the Local Group will have relatively small fields of view; thus, massive central galaxies will provide natural targets around which to search for dwarf satellite galaxies in deep pointings. We focus on MW-mass systems because lower-mass centrals host fewer satellites and because MW-mass systems are prevalent in the Local Volume (e.g., Mutlu-Pakdil et al. 2021; Carlsten et al. 2022) and nearby universe (e.g., Geha et al. 2017; Mao et al. 2021).
By combining our simulations with an expanded version of the galaxy–halo connection model and probabilistic framework from Nadler et al. (2019, 2020b, 2021c), we explore degeneracies between galaxy formation and DM effects on dwarf satellite populations and quantify signals of new physics that future galaxy surveys can uncover. We forecast that observations of all satellite galaxies around two MW-mass hosts (e.g., the MW and M31 or another nearby galaxy) can constrain a galaxy formation cutoff at peak virial halo masses of $10^{8}~{}M_{\mathrm{\odot}}$ and place stringent upper limits on cutoffs at lower halo masses. We show that the same data can constrain WDM models with masses of $\approx 10$ to $20~{}\mathrm{keV}$, significantly improving upon current constraints (Bullock & Boylan-Kolchin, 2017; Drlica-Wagner et al., 2019). Furthermore, we find that the SHMF can be measured in a model-independent fashion at peak virial masses between $10^{8}~{}M_{\mathrm{\odot}}$ to $10^{10}~{}M_{\mathrm{\odot}}$, paving the way for tests of any DM or early-universe physics that suppresses or enhances low-mass halo abundances relative to CDM.
@@
We adopt the same cosmological parameters used for our MW zoom-in simulations: $h=0.7$, $\Omega_{\rm m}=0.286$, $\Omega_{\Lambda}=0.714$, $\sigma_{8}=0.82$, and $n_{s}=0.96$ (Hinshaw et al., 2013). We quote (sub)halo virial masses using the Bryan & Norman (1998) virial overdensity, which corresponds to $\Delta_{\mathrm{vir}}\approx 99$ times the critical density of the universe at $z=0$ in this cosmology. Throughout, we refer to DM halos within a host’s virial radius as “subhalos”, and we denote galaxies that occupy a host’s subhalos as “satellites”. Furthermore, “log” always refers to the base-$10$ logarithm.
@@ 2.1 Cosmological Zoom-in Simulations
Our forecasts are based on subhalo populations from cosmological dark matter-only zoom-in simulations of MW-mass systems. Specifically, we use the “Milky Way-mass” suite from the Symphony zoom-in compilation (Mao et al., 2015; Nadler et al., 2023). This suite contains $45$ host halos in a virial mass range of $10^{12.09\pm 0.02}~{}M_{\mathrm{\odot}}$ and resolves subhalo abundances with present-day virial masses of $M_{\mathrm{vir}}>1.2\times 10^{8}~{}M_{\mathrm{\odot}}$. Subhalos below the convergence threshold derived in Nadler et al. (2023) contribute negligibly to the observable satellite populations in our forecasts and thus do not impact our results. We use the entire subhalo population within the virial radius of each MW-mass system—with no additional constraints imposed on e.g. subhalo distance from the host center, infall time, or size—to forward-model the entire observable satellite population of each host.
Our simulated MW-mass hosts have a range of secondary halo properties and inhabit a variety of cosmic environments (Fielder et al., 2019; Nadler et al., 2023). Thus, host-to-host scatter in subhalo populations, which results in subhalo abundance variations up to a factor of $\approx 2$ at fixed host halo mass (Mao et al., 2015), is naturally included in our predictions. However, we do not attempt to model the specific region of the nearby universe inhabited by the closest forty MW-mass systems, which provide natural observational targets for our forecasts. In particular, our simulations span a narrow range of host halo mass, whereas there is substantial variation in stellar mass and inferred host halo mass for nearby central galaxies. For example, the volume-limited sample of $31$ massive hosts within $12~{}\mathrm{Mpc}$ from Carlsten et al. (2022) spans central stellar masses of $9.9\lesssim\log(M_{*}/M_{\mathrm{\odot}})\lesssim 11.1$, and Danieli et al. (2023) models $27$ hosts from this population (excluding galaxy groups) with semi-analytic realizations of host halos with $10.5\lesssim\log(M_{\mathrm{host}}/M_{\mathrm{\odot}})\lesssim 13.3$. Our forecasts should therefore be viewed as illustrative estimates of the constraining power from future dwarf satellite galaxy surveys rather than specific predictions for upcoming observations. We discuss aspects of our simulations and galaxy–halo connection model that may be generalized to capture a broader range of host halo masses, secondary properties, and environments in Section 7.3.
@@ 2.2 Galaxy–Halo Connection Model
We use the CDM galaxy–halo connection model from (Nadler et al., 2019, 2020b) with a few minor modifications described below. This model associates satellite galaxies with subhalos using extrapolated abundance-matching relations to predict satellite luminosity and size. Each satellite’s absolute $V$-band magnitude, $M_{V}$, is predicted from its subhalo’s peak maximum circular velocity, $V_{\mathrm{peak}}$, using an extrapolated abundance-matching relation anchored to the GAMA luminosity function at $M_{r}=-13~{}\mathrm{mag}$ (Loveday et al., 2015), assuming a value for the faint-end luminosity function slope $\alpha$, and for the scatter $\sigma_{M}$ in absolute magnitude at fixed $V_{\mathrm{peak}}$. [^2] After performing abundance matching, we convert $r$-band to $V$-band magnitude via $M_{V}=M_{r}+0.2~{}\mathrm{mag}$ (Geha et al., 2017; Nadler et al., 2019). Both $\sigma_{M}$ and $\alpha$ are free parameters in our forecasts. Stellar mass–halo mass relation constraints derived from independent galaxy samples (e.g., Danieli et al. 2023) are consistent with the relation derived from GAMA, lending confidence to our approach.
@@
where $R_{\mathrm{vir,acc}}$ denotes a subhalo’s virial radius, measured when it accretes into its host’s virial radius; $R_{0}=10\ \mathrm{kpc}$ is a fixed normalization constant. The amplitude parameter $\mathcal{A}$ and power-law slope $n$ are free parameters in our analyses. After mean sizes are predicted for a given set of model parameters, each satellite’s half-light radius is drawn from a lognormal distribution with scatter $\sigma_{\log R}$, which is also a free parameter. A power-law galaxy–halo size relation as in Equation 1 has been shown to fit a variety of hydrodynamic simulation results reasonably well (e.g., Jiang et al. 2019), although some semi-analytic models (SAMs) predict a break in the size relation for ultra-faint dwarf galaxies, which we do not attempt to model (e.g., Kravtsov & Manwadkar 2022).
@@
where $p_{\mathrm{disrupt},0}$ is the fiducial subhalo disruption probability assigned by the Nadler et al. (2018) model, which depends on each subhalo’s mass and maximum circular velocity at infall, accretion time, and orbital properties. Each satellite is weighted by $1-p_{\mathrm{disrupt}}$ when number counts are computed from realizations of our model. In Equation 3, $\mathcal{B}$ is a free parameter that controls the strength of disruption. In particular, smaller (larger) values of $\mathcal{B}$ correspond to weaker (stronger) disruption due to baryons, and $\mathcal{B}=1$ corresponds to the fiducial amount of disruption measured in FIRE simulations. We adopt a wide prior on $\mathcal{B}$ because the efficiency of subhalo disruption due to central galaxies (e.g., Webb & Bovy 2020; Green et al. 2022) and the potential numerical origins of CDM subhalo disruption in general (van den Bosch et al., 2018; Green et al., 2021; Errani & Navarro, 2021; Benson & Du, 2022; Mansfield et al., 2023) are highly uncertain; our forecasts conservatively bracket these uncertainties.
For simplicity, we do not include orphan subhalos in our forecasts. Nadler et al. (2020b) showed, using the orphan model from Nadler et al. (2019), that orphans do not affect galaxy–halo connection constraints derived from current MW satellite observations in our modeling framework (however, see Bose et al. 2020). Because our main analyses do not consider fainter satellites (and thus lower-mass subhalos) than the Nadler et al. (2020b) analysis, we do not expect orphans to impact our results. Moreover, the convergence analyses in Nadler et al. (2023) indicate that our simulations accurately predict subhalo abundances above the fiducial resolution cuts described above. Relaxing this resolution cut would require modeling highly-stripped subhalos, which is difficult for standard halo finders (Mansfield et al., 2023) but will be an important area for future work.

### v2

@@ Abstract
The abundance of faint dwarf galaxies is determined by the underlying population of low-mass dark matter (DM) halos and the efficiency of galaxy formation in these systems. Here, we quantify potential galaxy formation and DM constraints from future dwarf satellite galaxy surveys. We generate satellite populations using a suite of Milky Way (MW)–mass cosmological zoom-in simulations and an empirical galaxy–halo connection model, and assess sensitivity to galaxy formation and DM signals when marginalizing over galaxy–halo connection uncertainties. We find that a survey of all satellites around one MW-mass host can constrain a galaxy formation cutoff at peak virial masses of $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$ at the $1\sigma$ level; however, a tail toward low $\mathcal{M}_{50}$ prevents a $2\sigma$ measurement. In this scenario, combining hosts with differing bright satellite abundances significantly reduces uncertainties on $\mathcal{M}_{50}$ at the $1\sigma$ level, but the $2\sigma$ tail toward low $\mathcal{M}_{50}$ persists. We project that observations of one (two) complete satellite populations can constrain warm DM models with $m_{\mathrm{WDM}}\approx 10~{}\mathrm{keV}$ ($20~{}\mathrm{keV}$). Subhalo mass function (SHMF) suppression can be constrained to $\approx 70\%$, $60\%$, and $50\%$ that in cold dark matter (CDM) at peak virial masses of $10^{8}$, $10^{9}$, and $10^{10}~{}M_{\mathrm{\odot}}$, respectively; SHMF enhancement constraints are weaker ($\approx 20$, $4$, and $2$ times that in CDM, respectively) due to galaxy–halo connection degeneracies. These results motivate searches for faint dwarf galaxies beyond the MW and indicate that ongoing missions like Euclid and upcoming facilities including the Vera C. Rubin Observatory and Nancy Grace Roman Space Telescope will probe new galaxy formation and DM physics.
@@ 1 Introduction
The faintest dwarf galaxies inhabit the smallest dark matter halos that can form galaxies. These systems are sensitive probes of galaxy formation models and simultaneously test dark matter (DM) particle physics (e.g., Bullock & Boylan-Kolchin 2017; Gluscevic et al. 2019; Simon 2019; Nadler 2021; Sales et al. 2022), including DM free-streaming, interactions, and decays, wave interference of ultralight DM, and nonstandard DM production mechanisms (e.g., see Bechtol et al. 2022 for an overview).
@@
To disentangle the information dwarf galaxy surveys can offer, it is crucial to understand the interplay between galaxy formation and DM physics in low-mass halos. For example, reionization inhibits star formation in halos with peak masses below $\approx 10^{8}$ to $10^{9}~{}M_{\mathrm{\odot}}$ because of their low virial temperatures and limited gas reservoirs (e.g., Benitez-Llambay & Frenk 2020; Munshi et al. 2021; Kravtsov & Manwadkar 2022; Ahvazi et al. 2024), suppressing the population of dwarf galaxies that occupy these systems. Observations of the faintest dwarf galaxies can shed light on these processes and can also reveal deviations in underlying halo populations from cold dark matter (CDM), which would signal new DM physics on small scales. Moreover, a measurement of the galaxy formation cutoff is necessary to infer the existence of “dark” (galaxy–free) halos—which remain a key, untested predicted of many DM models—in combination with gravitational probes of even smaller halos (e.g., Gilman et al. 2020a; Banik et al. 2021b; Nadler et al. 2021b).
@@
Current constraints on the galaxy formation cutoff are limited by the small number of known ultrafaint dwarf galaxies, which have almost exclusively been detected as satellites of the Milky Way (MW; e.g., Drlica-Wagner et al. 2020). Additionally, this measurement is plagued by large theoretical uncertainties on the faint end of the galaxy–halo connection (e.g., see Wechsler & Tinker 2018 for a review). Recently, Nadler et al. (2020b) analyzed the full-sky MW satellite population detected by the Dark Energy Survey (DES; Abbott et al. 2018) and Pan-STARRS1 (PS1; Chambers et al. 2016) to place a $95\%$ confidence upper limit of $8.5\times 10^{7}~{}M_{\mathrm{\odot}}$ on the peak virial mass at which $50\%$ of halos host galaxies, hereafter $\mathcal{M}_{50}$. [^1] We define “galaxies” as dark matter-dominated systems with absolute $V$-band magnitudes $M_{V}<0~{}\rm{mag}$, which are composed of $\gtrsim 100$ stars in our fiducial models, on average, following Nadler et al. (2020b). This constraint accounts for observational selection effects, marginalizes over the relevant theoretical uncertainties, and implies stringent limits on a variety of DM scenarios that suppress the abundance of halos with peak masses of $\approx 10^{8}~{}M_{\mathrm{\odot}}$ (e.g., Nadler et al. 2021c; Newton et al. 2021; Dekker et al. 2022). However, current data do not permit a measurement of the galaxy formation cutoff because only tens of ultrafaints have been detected; this number is too small to rule out statistical fluctuations that mimic a cutoff or to overcome theoretical uncertainties on the predicted abundance of MW satellite galaxies.
@@
Upcoming observational facilities will substantially improve this situation by expanding our census of faint dwarf galaxies. For example, the Vera C. Rubin Observatory (Rubin; Ivezić et al. 2008) is expected to detect much of the remaining MW satellite population, as well as many resolved dwarf galaxies out to $\approx 5~{}\mathrm{Mpc}$ (Drlica-Wagner et al., 2019; Mutlu-Pakdil et al., 2021); comparable sensitivity has been projected for the Chinese Space Station Telescope (Qu et al., 2023). The Nancy Grace Roman Space Telescope (Roman; Spergel et al. 2015) will also be a powerful dwarf galaxy discovery machine (Gezari et al., 2022), with the ability to detect resolved stars out to even larger distances than Rubin. The Euclid Dark Energy Mission (Euclid Collaboration et al. 2022; launched 2023 July 1) is expected to achieve similar sensitivity (Racca et al., 2016). Meanwhile, despite its small field of view, JWST can resolve stars out to much larger distances than Roman or Euclid and thus may play an important role in detecting faint dwarf galaxies throughout the Local Volume and beyond (e.g., Conroy et al. 2023). Looking further ahead, next-generation optical/infrared space telescopes such as the Habitable Worlds Observatory, which is the top priority of the Astro2020 Decadal Survey, will revolutionize the search for dwarf galaxies at cosmic distances (e.g., see The LUVOIR Team 2019 for estimates of future large space telescopes’ detection sensitivities).
@@
To prepare for these advances, we study measurements of galaxy formation and DM physics in the context of future dwarf galaxy surveys. In particular, we use a suite of MW-mass cosmological zoom-in simulations (Nadler et al., 2023) to assess the sensitivity of satellite galaxy populations to a reionization-like galaxy formation cutoff, a suppression of low-mass halo abundances characteristic of warm dark matter (WDM), and the amplitude of the subhalo mass function (SHMF). We focus on subhalos and satellite galaxies (rather than isolated halos and field galaxies) because, as discussed above, future observational facilities capable of resolving ultrafaint dwarfs beyond the Local Group will have relatively small fields of view; thus, massive central galaxies will provide natural targets around which to search for dwarf satellite galaxies in deep pointings. We focus on MW-mass systems because lower-mass centrals host fewer satellites and because MW-mass systems are prevalent in the Local Volume (e.g., Mutlu-Pakdil et al. 2021; Carlsten et al. 2022) and nearby Universe (e.g., Geha et al. 2017; Mao et al. 2021).
[…]
@@
We adopt the same cosmological parameters used for our MW zoom-in simulations: $h=0.7$, $\Omega_{\rm m}=0.286$, $\Omega_{\Lambda}=0.714$, $\sigma_{8}=0.82$, and $n_{s}=0.96$ (Hinshaw et al., 2013). We quote (sub)halo virial masses using the Bryan & Norman (1998) virial overdensity, which corresponds to $\Delta_{\mathrm{vir}}\approx 99$ times the critical density of the Universe at $z=0$ in this cosmology. Throughout, we refer to DM halos within a host’s virial radius as “subhalos,” and we denote galaxies that occupy a host’s subhalos as “satellites.” Furthermore, “log” always refers to the base-$10$ logarithm.
@@ 2.1 Cosmological Zoom-in Simulations
Our forecasts are based on subhalo populations from cosmological dark matter-only zoom-in simulations of MW-mass systems. Specifically, we use the “Milky Way-mass” suite from the Symphony zoom-in compilation (Mao et al., 2015; Nadler et al., 2023). This suite contains $45$ host halos in a virial mass range of $10^{12.09\pm 0.02}~{}M_{\mathrm{\odot}}$ and accurately resolves the abundance of subhalos with present-day virial masses of $M_{\mathrm{vir}}>1.2\times 10^{8}~{}M_{\mathrm{\odot}}$ (Nadler et al., 2023). Subhalos below this convergence threshold contribute negligibly to the observable satellite populations in our main forecasts, which are restricted to dwarf galaxies with $M_{V}<0~{}\mathrm{mag}$ that occupy subhalos with $M_{\mathrm{vir}}>10^{8}~{}M_{\mathrm{\odot}}$, on average (Nadler et al., 2020b). Thus, poorly resolved subhalos do not impact our results. We use the entire subhalo population within the virial radius of each MW-mass system—with no additional constraints imposed on e.g. subhalo distance from the host center, infall time, or size—to forward-model the entire observable satellite population of each host.
Our simulated MW-mass hosts have a range of secondary halo properties and inhabit a variety of cosmic environments (Fielder et al., 2019; Nadler et al., 2023). Thus, host-to-host scatter in subhalo populations, which results in subhalo abundance variations up to a factor of $\approx 2$ at fixed host halo mass (Mao et al., 2015), is naturally included in our predictions. However, we do not attempt to model the specific region of the nearby Universe inhabited by the closest $40$ MW-mass systems, which provide natural observational targets for our forecasts. In particular, our simulations span a narrow range of host halo mass, whereas there is substantial variation in stellar mass and inferred host halo mass for nearby central galaxies. For example, the volume-limited sample of $31$ massive hosts within $12~{}\mathrm{Mpc}$ from Carlsten et al. (2022) spans central stellar masses of $9.9\lesssim\log(M_{*}/M_{\mathrm{\odot}})\lesssim 11.1$, and Danieli et al. (2023) models $27$ hosts from this population (excluding galaxy groups) with semianalytic realizations of host halos with $10.5\lesssim\log(M_{\mathrm{host}}/M_{\mathrm{\odot}})\lesssim 13.3$. Our forecasts should therefore be viewed as illustrative estimates of the constraining power from future dwarf satellite galaxy surveys rather than specific predictions for upcoming observations. We discuss aspects of our simulations and galaxy–halo connection model that may be generalized to capture a broader range of host halo masses, secondary properties, and environments in Section 7.3.
@@ 2.2 Galaxy–Halo Connection Model
We use the CDM galaxy–halo connection model from Nadler et al. (2019b, 2020b) with a few minor modifications described below. This model associates satellite galaxies with subhalos using extrapolated abundance-matching relations to predict satellite luminosity and size. Each satellite’s absolute $V$-band magnitude, $M_{V}$, is predicted from its subhalo’s peak maximum circular velocity, $V_{\mathrm{peak}}$, using an extrapolated abundance-matching relation anchored to the GAMA luminosity function at $M_{r}=-13~{}\mathrm{mag}$ (Loveday et al., 2015), assuming a value for the faint-end luminosity function slope $\alpha$, and for the scatter $\sigma_{M}$ in absolute magnitude at fixed $V_{\mathrm{peak}}$. [^2] After performing abundance matching, we convert $r$-band to $V$-band magnitude via $M_{V}=M_{r}+0.2~{}\mathrm{mag}$ (Geha et al., 2017; Nadler et al., 2019b). Both $\sigma_{M}$ and $\alpha$ are free parameters in our forecasts. Stellar mass–halo mass relation constraints derived from independent galaxy samples (e.g., Danieli et al. 2023) are consistent with the relation derived from GAMA, lending confidence to our approach.
@@
where $R_{\mathrm{vir,acc}}$ denotes a subhalo’s virial radius, measured when it accretes into its host’s virial radius; $R_{0}=10\ \mathrm{kpc}$ is a fixed normalization constant. The amplitude parameter $\mathcal{A}$ and power-law slope $n$ are free parameters in our analyses. After mean sizes are predicted for a given set of model parameters, each satellite’s half-light radius is drawn from a lognormal distribution with scatter $\sigma_{\log R}$, which is also a free parameter. A power-law galaxy–halo size relation as in Equation 1 has been shown to fit a variety of hydrodynamic simulation results reasonably well (e.g., Jiang et al. 2019), although some semianalytic models (SAMs) predict a break in the size relation for ultrafaint dwarf galaxies, which we do not attempt to model (e.g., Kravtsov & Manwadkar 2022).
@@
where $p_{\mathrm{disrupt},0}$ is the fiducial subhalo disruption probability assigned by the Nadler et al. (2018) model, which depends on each subhalo’s mass and maximum circular velocity at infall, accretion time, and orbital properties. Note that we do not modify $p_{\mathrm{disrupt},0}$ in our beyond-CDM scenarios because it mainly depends on a given subhalo’s accretion time and orbital properties (Garrison-Kimmel et al., 2017b; Nadler et al., 2018), which do not significantly differ between CDM and models like WDM (e.g., Knebe et al. 2008).
To implement the disruption model in our analysis, each satellite is weighted by $1-p_{\mathrm{disrupt}}$ when number counts are computed from realizations of our model. In Equation 3, $\mathcal{B}$ is a free parameter that controls the strength of disruption. In particular, smaller (larger) values of $\mathcal{B}$ correspond to weaker (stronger) disruption due to baryons, and $\mathcal{B}=1$ corresponds to the fiducial amount of disruption measured in FIRE simulations. We adopt a wide prior on $\mathcal{B}$ because the efficiency of subhalo disruption due to central galaxies (e.g., Webb & Bovy 2020; Green et al. 2022) and the potential numerical origins of CDM subhalo disruption in general (van den Bosch et al., 2018; Errani & Navarro, 2021; Green et al., 2021; Benson & Du, 2022; Mansfield et al., 2023) are highly uncertain; our forecasts conservatively bracket these uncertainties.
For simplicity, we do not include orphan subhalos in our forecasts. Nadler et al. (2020b) showed, using the orphan model from Nadler et al. (2019b), that orphans do not affect galaxy–halo connection constraints derived from current MW satellite observations in our modeling framework (however, see Bose et al. 2020). Because our main analyses do not consider fainter satellites (and thus lower-mass subhalos) than the Nadler et al. (2020b) analysis, we do not expect orphans to impact our results. Moreover, the convergence analyses in Nadler et al. (2023) indicate that our simulations accurately predict subhalo abundances above the fiducial resolution cuts described above. Modeling dwarf galaxies that occupy halos with $M_{\mathrm{vir}}\lesssim 10^{8}~{}M_{\mathrm{\odot}}$ would require modeling highly stripped subhalos, which is difficult for standard halo finders (Mansfield et al., 2023) but will be an important area for future work.